# Quantum framework for Reinforcement Learning: integrating Markov Decision Process, quantum arithmetic, and trajectory search

### v2

@@ V.2 Demonstration of agent-environment quantum interactions over multiple time steps
In this section, we extend the QMDP model to simulate multiple agent-environment interactions over 3 consecutive time steps, focusing on how the state evolves based on repeated interactions of the agent and the environment. Figure 7 illustrates the QMDP circuit for this 3-time step process, where each time step (labeled $t0,t1$, and $t2$) represents a unique interaction between the agent and the environment. To model the state transitions between time steps, controlled-Not gates are employed at each step. These gates update the state qubits of the following time step conditionally, depending on the evolution of the new state in response to the prior interaction of the agent with the environment. This configuration allows the QMDP circuit to evaluate not just individual state transitions, but also the cumulative effect of the agent’s interactions over 3 time steps, effectively capturing the rewards across multiple agent and environment interactions.
Figure 7: Quantum circuit implementation of agent-environment interactions across 3 time steps (t = 0, 1, 2). Each colored block represents a single time step, with qubits for states (qState), actions (qAction), next states (qNextState), and rewards (qReward). Controlled-Not gates represent state transitions between the interactions.
@@ V.3 Demonstration of quantum trajectories search
In this section, we present the results of implementing Grover’s search algorithm to identify optimal quantum trajectories within the quantum reinforcement learning (QRL) framework outlined in the method section. The objective of this demonstration is to efficiently search through the set of possible quantum trajectories generated from agent-environment interactions to find those that maximize the overall return, thereby guiding the discovery of optimal policies for the QRL agent. Each trajectory is composed of a sequence of quantum states, actions, next state, rewards, and a return value, and Grover’s search algorithm is employed to efficiently identify those trajectories with the highest return from a large set of possible agent-environment interactions.
@@
In Table 2, each quantum sample for a given time step is represented by a structured sequence of qubits. Starting from the least significant bits, the first two qubits represent the current state, followed by the action qubit, then the next two qubits indicate the next state, and finally, the two most significant bits represent the reward. These qubit representations are described in Table 1.
@@
To compare these quantum results with a classical RL approach, we conducted an analysis using Q-learning on an equivalent MDP dynamic. For this comparison, we set a discount factor of 1 to allow us to evaluate the consistency and effectiveness of the trajectories identified by Grover’s search with those found through classical Q-learning. The Q-table learned from the Q-learning algorithm is described in Table 3, which presents the Q-values generated by the algorithm for an environment with four states ($s_{0},s_{1},s_{2},s_{3}$) and two actions ($a_{0},a_{1}$). Each cell in the table contains the Q-value for a state-action pair, which represents the expected cumulative reward when taking a respective action in that given state. For each state, the action with the highest Q-value is highlighted in yellow, indicating the optimal action to maximize expected rewards. In state $s_{0}$, action $a_{0}$ has the highest Q-value (38.02), making it optimal. In states $s_{1}$, $s_{2}$, and $s_{3}$, action $a_{1}$ is optimal, with Q-values of 47.89, 54.28, and 59.54, respectively.
@@
After generating the Q-table, we executed the Q-learning algorithm over 100 trials, each running for 3 time steps, to identify and evaluate potential trajectories. This process resulted in four unique trajectories, each with an associated total reward, as shown in Fig. 11. Figure 11 provides a comparison of total rewards for these four trajectories, beginning at state $s_{0}$ and terminating at state $s_{3}$. The y-axis represents the total reward, while the x-axis lists the different trajectories. Each bar represents a distinct trajectory, labeled $T1,T2,T3,$ and $T4$, with the total reward achieved by each trajectory shown above the corresponding bar. Among these, trajectory $T1$ achieves the highest total reward of $8$, identifying it as the optimal path. In $T1$, the agent starts at $s_{0}$, takes action $a_{0}$ to reach $s_{2}$ and receive $r_{2}$ which is represented as “(0,0,2,2)”, then takes action $a_{1}$ to move to $s_{3}$ with $r_{3}$ described as “(2,1,3,3)”, and finally takes $a_{1}$ again to remain in $s_{3}$ representing “(3,1,3,3)”. This sequence results in the highest total reward among all trajectories, indicating that it is the most rewarding path available within the constraints.
@@ V.4 Advantages of quantum RL over classical approaches
By representing both the agent and environment as quantum components and leveraging quantum superposition, the quantum model enables the simultaneous exploration of multiple trajectories. This parallelism allows the quantum agent to evaluate numerous interaction sequences in a single execution, significantly reducing the number of interactions required to explore the environment. As a result, the quantum approach achieves much greater sample efficiency compared to classical methods, which typically process one interaction sequence at a time and require many iterations to converge on an optimal policy.
@@ VII Conclusion
Future research could be extended by exploring more complex MDPs with larger state and action spaces, addressing scalability challenges by optimizing qubit usage. Investigating alternative quantum search algorithms beyond Grover’s could further enhance efficiency for various decision-making scenarios. These advancements could expand the applicability of quantum MDPs and drive progress in quantum reinforcement learning for complex, resource-intensive tasks.
@@ Acknowledgements.
This work was supported by JST COI-NEXT Grant Number JPMJPF2221, JSPS KAKENHI Grant Number JP24K20843, and the Ministry of Education, Culture, Sports, Science, and Technology (MEXT), Japan. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.

### v3

@@ V.2 Demonstration of agent-environment quantum interactions over multiple time steps
In this section, we extend the QMDP model to simulate multiple agent-environment interactions over 3 consecutive time steps, focusing on how the state evolves based on repeated interactions of the agent and the environment. Figure 7 illustrates the QMDP circuit for this 3-time step process, where each time step (labeled $t0,t1$, and $t2$) represents a unique interaction between the agent and the environment. To model the state transitions between time steps, CNOT gates are employed at each step. These gates update the state qubits of the following time step conditionally, depending on the evolution of the new state in response to the prior interaction of the agent with the environment. This configuration allows the QMDP circuit to evaluate not just individual state transitions, but also the cumulative effect of the agent’s interactions over 3 time steps, effectively capturing the rewards across multiple agent and environment interactions.
Figure 7: Quantum circuit implementation of agent-environment interactions across 3 time steps (t = 0, 1, 2). Each colored block represents a single time step, with qubits for states (qState), actions (qAction), next states (qNextState), and rewards (qReward). CNOT gates represent state transitions between the interactions.
@@ V.3 Demonstration of quantum trajectories search
In this section, we present the results of implementing Grover’s search algorithm to identify optimal quantum trajectories within the quantum reinforcement learning (QRL) framework outlined in the method section. The objective of this demonstration is to efficiently search through the set of possible quantum trajectories generated from agent-environment interactions to find those that maximize the overall return, thereby guiding the discovery of optimal policies for the QRL agent. Each trajectory is composed of a sequence of states, actions, next states, rewards, and a return value, and Grover’s search algorithm is employed to efficiently identify those trajectories with the highest return from a large set of possible agent-environment interactions.
@@
In Table 2, each quantum sample for a given time step is represented by a structured sequence of qubits. The qubit representation for each time step is detailed in Table 1.
@@
To compare these quantum results with a classical RL approach, we conducted an analysis using Q-learning on an equivalent MDP dynamic. For this comparison, we set a discount factor of 1 to allow us to evaluate the consistency and effectiveness of the trajectories identified by Grover’s search with those found through classical Q-learning. In an initial step of the Q-learning process, the Q-table was set to zero and updated through iterative runs of the Q-learning algorithm until it converged toward the optimal policy. Table 3 presents the resulting Q-table, which contains the Q-values learned for an environment with four states ($s_{0},s_{1},s_{2},s_{3}$) and two actions ($a_{0},a_{1}$). Each cell in the table contains the Q-value for a state-action pair, which represents the expected cumulative reward when taking a respective action in that given state. For each state, the action with the highest Q-value is highlighted in yellow, indicating the optimal action to maximize expected rewards. In state $s_{0}$, action $a_{0}$ has the highest Q-value (38.02), making it optimal. In states $s_{1}$, $s_{2}$, and $s_{3}$, action $a_{1}$ is optimal, with Q-values of 47.89, 54.28, and 59.54, respectively.
@@
After generating the Q-table, we used the learned Q-values to evaluate potential trajectories by running the Q-learning algorithm over 100 trials, each consisting of 3 time steps. This process resulted in four unique trajectories, each with an associated total reward, as shown in Fig. 11. Figure 11 provides a comparison of total rewards for these four trajectories, beginning at state $s_{0}$ and terminating at state $s_{3}$. The y-axis represents the total reward, while the x-axis lists the different trajectories. Each bar represents a distinct trajectory, labeled $T1,T2,T3,$ and $T4$, with the total reward achieved by each trajectory shown above the corresponding bar. Among these, trajectory $T1$ achieves the highest total reward of $8$, identifying it as the optimal path. In $T1$, the agent starts at $s_{0}$, takes action $a_{0}$ to reach $s_{2}$ and receive $r_{2}$ which is represented as “(0,0,2,2)”, then takes action $a_{1}$ to move to $s_{3}$ with $r_{3}$ described as “(2,1,3,3)”, and finally takes $a_{1}$ again to remain in $s_{3}$ representing “(3,1,3,3)”. This sequence results in the highest total reward among all trajectories, indicating that it is the most rewarding path available within the constraints.
@@ V.4 Advantages of quantum RL over classical approaches
By representing both the agent and environment as quantum components and leveraging quantum computations, our quantum framework enables the quantum agent to evaluate numerous interaction sequences in a single execution, significantly reducing the number of interactions required to explore the environment. As a result, the quantum approach achieves much greater sample efficiency compared to classical methods, which typically process one interaction sequence at a time and require many iterations to converge on an optimal policy.
@@ VII Conclusion
Future research could be extended by exploring more complex MDPs with larger state and action spaces, addressing scalability challenges by optimizing qubit usage. In our current implementation of Grover’s search, we assume prior knowledge of the maximum return and construct the quantum oracle accordingly. A promising direction for future work is to develop quantum approaches capable of searching for unknown optimal returns without this prior knowledge. Additionally, since qubit usage increases significantly when running over multiple time steps, investigating techniques for reusing qubits could help mitigate resource demands. Exploring alternative quantum search algorithms beyond Grover’s could further enhance efficiency in various decision-making scenarios. These advancements could expand the applicability of quantum MDPs and drive progress in quantum reinforcement learning for complex, resource-intensive tasks.
@@ Acknowledgements.
This work was supported by the Center of Innovations for Sustainable Quantum AI (JST Grant Number JPMJPF2221), JSPS KAKENHI Grant Number JP24K20843, and the Ministry of Education, Culture, Sports, Science and Technology (MEXT) of Japan. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.